A student finishes the first half of an exam in 2/3 the time it takes him to finish the second half. If the entire exam takes him an hour, how many minutes does he spend on the first of the exam?
Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

Let x=time to finish 2nd half of the exam.
2x/3 + x = 60
2x+3x=180 multiply each side by 3 to eliminate fraction.
5x=180
x=36
2*36/3=24 min to finish 1st half of the exam.
.
Ed

Answer by stanbon(75887)

A student finishes the first half of an exam in 2/3 the time it takes him to finish the second half. If the entire exam takes him an hour, how many minutes does he spend on the first of the exam?
------------------------------------------------
Let the time for the 2nd half be "x".
Then the 1st half takes (2/3)x
----------------------------------
Equation:
x + (2/3)x = 60 minutes
(5/3)x = 60
5x = 180
x = 36 minutes
(2/3)x = (2/3)36 = 24 minutes (time spent on the 1st half of the exam)
==========================================
Cheers,
Stan H.